OpenStudy (anonymous):

Give the equation of a line that has a slope of 4 and passes through the point (3, –2).

OpenStudy (anonymous):

A. y = 4x – 14 B. y = –2x + 11 C. y = 3x – 14 D. y = –2x – 9

OpenStudy (ryan123345):